OpenStudy (anonymous):

integral of (7x/(x^2-4))dx

OpenStudy (mimi_x3):

\[\large \int\limits\frac{7x}{x^{2}-4} dx\] let u= x^2 - 4 => du/ 2x \[\large \int\limits\frac{7x}{u} *\frac{du}{2x} \] \[\large \frac{7}{2} \int\limits\frac{1}{u} du\] Then integrate it..